# InfinityStar: Unified Spacetime AutoRegressive Modeling for Visual Generation

### v1

@@ 4.3 Text-to-Video Generation
In the lower part of Fig.5, we present the generated videos of InfinityStar regarding user prompts. The generated videos successfully capture the semantic information in user prompts while maintaining high aesthetics and visual quality. Especially for the second example in Fig.5, the generated video accurately restores the delicate movements of the characters flipping through sketchbooks, talking while pointing to different parts of the drawings. In Tab.2, we compare InfinityStar with leading diffusion and autoregressive approaches on VBench—a comprehensive video benchmark spanning 16 evaluation dimensions. Our model achieves an overall score of 83.74, outperforming all open-source autoregressive baselines by a substantial margin. Moreover, InfinityStar surpasses diffusion-based competitors such as OpenSora[50], CogVideoX[40], and HunyuanVideo[19]. These results demonstrate that, through its novel spacetime autoregressive design, InfinityStar not only pushes the capabilities of discrete autoregressive video models but also attains performance on par with—and in some cases superior to—state-of-the-art diffusion methods.
@@ 5.1 Model Design
Beyond spacetime sparse attention, we introduce the novel Semantic-Detail conditions to control video synthesizing when interacting with users as illustrated in Fig.10. Specifically, we extract features $F_{i-1}\in T\times H\times W$ from the preceding clip $V^{1}_{i-1}$ using the visual tokenizer. The features $F_{i-1}$ are referred to detail features since they are full-scale and contain rich visual details. It is difficult to extract semantic information from $F_{i-1}$ because it is not adequately compressed. Besides, there are too many tokens in $F_{i-1}$, which significantly slows down the interactive inference speed. Borrow ideas from FramePack [46], we downsample $F_{i-1}$ to $F_{i-1}^{sem}\in T\times h\times w$ spatially to reduce excessive condition tokens. The semantic conditions $F_{i-1}^{sem}$ are employed to enable semantic consistency between clips. Apart from $F_{i-1}^{sem}$, we slice the last $K$ frames from $F_{i-1}$ instead of the whole as the detail conditions $F_{i-1}^{det}\in K\times H\times W$. In this way, we ensure consistency in both semantics and details while significantly compressing the number of condition tokens.
@@ 5.2 Dataset
We curate the interactive generation data from the pretraining dataset and other sources. In particular, we select videos longer than 7 seconds from the pretraining data, resulting in a total of 7M videos. Subsequently, we decompose long videos into chunks, split the chunks into clips, and generate captions at the clip level using the Tarsier2 [42] model. It is worth noting that here we adopt an LLM to remove the content that had already appeared in $V^{0}_{i}$’s caption from $V^{1}_{i}$’s caption, and ensure that $t^{1}_{i}$ only describes changes compared to $t^{0}_{i}$. The instructions used to query an LLM are presented in the Appendix C.4. In this way, we align with the instructions users provide during interactive generation.
@@ Appendix A Spacetime Autogressive Modeling
Spacetime RoPE. We introduce spacetime rotary position embeddings (Spacetime RoPE) tailored for InfinityStar. This is achieved by decomposing original rotary embeddings[25] into four components: scale, time, height, and width. As shown in Fig.14, the scale ID serves as a counter of scales up to now. The temporal ID remains zero for tokens in the image pyramid. For tokens in video pyramids, it increments as the frame grows. Distinct IDs are assigned to height and width components based on the token’s position in the image or video. Spacetime RoPE enhances the modeling of complex positional information for tokens in image and video pyramids.
Spacetime Autoregressive Transformer with Bitwise Self-Correction. To alleviate the train-test discrepancies of teacher-forcing training, we adopt bitwise self-correction mechanism proposed by Infinity[15]. Specifically, during training, some of the input tokens are randomly flipped to simulate the prediction error during the inference phase. Besides, the target labels are also recomputed to match the perturbed inputs. Moreover, when predicting the token distribution, the traditional index-wise classifier is replaced by a bitwise classifier. The bitwise classifier predicts $d$ bits instead of $2^{d}$ indices, significantly reducing the memory costs and difficulties in optimization. Algorithm 1 shows the detailed procedure of Spacetime Pyramid Encoding with Bitwise Self-Correction.
@@ Appendix B Infrastructure and Data
Firstly, we utilize FlexAttention to implement various attention mechanisms. With our proposed Spacetime Sparse Attention, we achieve more than a $2\times$ acceleration in training speed. Secondly, we adopt fully sharded data parallelism (FSDP) [48] to partition parameters, gradients, and optimizer states across GPU ranks. Thirdly, we adopt a fine-grained activation checkpointing strategy to reduce the overhead of vRAM and data transfer, making the parallelization more efficient. Last but not least, sequence parallelism further partitions long sequences into multiple chunks and then exploits ring self-attention for each chunk, making it feasible to train 720p videos with 200K sequence length.
Visual Captioning. Detailed visual captioning is crucial for enabling the model to accurately generate images and videos. For images, we use InternVL2.0[8] to produce dense descriptions for each sample. For video clips, we obtain overall video descriptions using Tarsier2[42]. Notably, Tarsier2 inherently captures camera motion types (e.g., zoom, pan right), eliminating the need for a separate prediction model. This simplifies the pipeline and improves efficiency.
Data Pipeline. Obtaining a high-quality image and video dataset requires a complex processing pipeline. Specifically for video, we follow video processing pipelines[6] to preprocess videos into high-quality training clips through OCR filtering, video clip extraction, visual aesthetic filtering, and motion filtering, etc.

### v2

@@ 4.3 Text-to-Video Generation
In the lower part of Fig.5, we present the generated videos of InfinityStar regarding user prompts. The generated videos successfully capture the semantic information in user prompts while maintaining high aesthetics and visual quality. Especially for the second example in Fig.5, the generated video accurately restores the delicate movements of the characters flipping through sketchbooks, talking while pointing to different parts of the drawings. In Tab.2, we compare InfinityStar with leading diffusion and autoregressive approaches on VBench—a comprehensive video benchmark spanning 16 evaluation dimensions. Our model achieves an overall score of 83.74, outperforming all open-source autoregressive baselines by a substantial margin. Moreover, InfinityStar surpasses diffusion-based competitors such as OpenSora[opensora], CogVideoX[cogvideox], and HunyuanVideo[Hunyuanvideo]. These results demonstrate that, through its novel spacetime autoregressive design, InfinityStar not only pushes the capabilities of discrete autoregressive video models but also attains performance on par with—and in some cases superior to—state-of-the-art diffusion methods.
@@ 5.1 Model Design
Beyond spacetime sparse attention, we introduce the novel Semantic-Detail conditions to control video synthesizing when interacting with users as illustrated in Fig.10. Specifically, we extract features $F_{i-1}\in T\times H\times W$ from the preceding clip $V^{1}_{i-1}$ using the visual tokenizer. The features $F_{i-1}$ are referred to detail features since they are full-scale and contain rich visual details. It is difficult to extract semantic information from $F_{i-1}$ because it is not adequately compressed. Besides, there are too many tokens in $F_{i-1}$, which significantly slows down the interactive inference speed. Borrow ideas from FramePack [framepack], we downsample $F_{i-1}$ to $F_{i-1}^{sem}\in T\times h\times w$ spatially to reduce excessive condition tokens. The semantic conditions $F_{i-1}^{sem}$ are employed to enable semantic consistency between clips. Apart from $F_{i-1}^{sem}$, we slice the last $K$ frames from $F_{i-1}$ instead of the whole as the detail conditions $F_{i-1}^{det}\in K\times H\times W$. In this way, we ensure consistency in both semantics and details while significantly compressing the number of condition tokens.
@@ 5.2 Dataset
We curate the interactive generation data from the pretraining dataset and other sources. In particular, we select videos longer than 7 seconds from the pretraining data, resulting in a total of 7M videos. Subsequently, we decompose long videos into chunks, split the chunks into clips, and generate captions at the clip level using the Tarsier2 [tarsier2] model. It is worth noting that here we adopt an LLM to remove the content that had already appeared in $V^{0}_{i}$’s caption from $V^{1}_{i}$’s caption, and ensure that $t^{1}_{i}$ only describes changes compared to $t^{0}_{i}$. The instructions used to query an LLM are presented in the Appendix C.4. In this way, we align with the instructions users provide during interactive generation.
@@ Appendix A Spacetime Autogressive Modeling
Spacetime RoPE. We introduce spacetime rotary position embeddings (Spacetime RoPE) tailored for InfinityStar. This is achieved by decomposing original rotary embeddings[ROPE] into four components: scale, time, height, and width. As shown in Fig.14, the scale ID serves as a counter of scales up to now. The temporal ID remains zero for tokens in the image pyramid. For tokens in video pyramids, it increments as the frame grows. Distinct IDs are assigned to height and width components based on the token’s position in the image or video. Spacetime RoPE enhances the modeling of complex positional information for tokens in image and video pyramids.
Spacetime Autoregressive Transformer with Bitwise Self-Correction. To alleviate the train-test discrepancies of teacher-forcing training, we adopt bitwise self-correction mechanism proposed by Infinity[Infinity]. Specifically, during training, some of the input tokens are randomly flipped to simulate the prediction error during the inference phase. Besides, the target labels are also recomputed to match the perturbed inputs. Moreover, when predicting the token distribution, the traditional index-wise classifier is replaced by a bitwise classifier. The bitwise classifier predicts $d$ bits instead of $2^{d}$ indices, significantly reducing the memory costs and difficulties in optimization. Algorithm 1 shows the detailed procedure of Spacetime Pyramid Encoding with Bitwise Self-Correction.
@@ Appendix B Infrastructure and Data
Firstly, we utilize FlexAttention to implement various attention mechanisms. With our proposed Spacetime Sparse Attention, we achieve more than a $2\times$ acceleration in training speed. Secondly, we adopt fully sharded data parallelism (FSDP) [fsdp] to partition parameters, gradients, and optimizer states across GPU ranks. Thirdly, we adopt a fine-grained activation checkpointing strategy to reduce the overhead of vRAM and data transfer, making the parallelization more efficient. Last but not least, sequence parallelism further partitions long sequences into multiple chunks and then exploits ring self-attention for each chunk, making it feasible to train 720p videos with 200K sequence length.
Visual Captioning. Detailed visual captioning is crucial for enabling the model to accurately generate images and videos. For images, we use InternVL2.0[chen2023internvl] to produce dense descriptions for each sample. For video clips, we obtain overall video descriptions using Tarsier2[tarsier2]. Notably, Tarsier2 inherently captures camera motion types (e.g., zoom, pan right), eliminating the need for a separate prediction model. This simplifies the pipeline and improves efficiency.
Data Pipeline. Obtaining a high-quality image and video dataset requires a complex processing pipeline. Specifically for video, we follow video processing pipelines[goku] to preprocess videos into high-quality training clips through OCR filtering, video clip extraction, visual aesthetic filtering, and motion filtering, etc.